OpenStudy (anonymous):

A game is played using one die. If the die is rolled and shows a 2, the player wins $7. If the die shows any number other than 2, the player wins nothing. If there is a charge of $1 to play the game, what is the game's expected value? Round to the nearest cent.

OpenStudy (anonymous):

is there are 6 sides on a die, and he can only get a 2. the chances of him winning are 1/6

OpenStudy (anonymous):

but what about the game, what is the chance of the game winning?

OpenStudy (anonymous):

or in other words what are the chances of the person losing?

OpenStudy (amistre64):

expected value is determined by the sum of xP(x)

OpenStudy (anonymous):

Number Outcome probability 1 $0 x 1/6 = $0 2 $7 x 1/6 = $1.166... 3 $0 x 1/6 = $0 etc., etc., Total : $1.166 Expected value = 1.166, but remember if there is a $1 charge, you will only get $0.166 (or $0.17c rounded to the nearest cent).